A Bibliography of Publications of Beny Neta

Beny Neta Naval Postgraduate School

Department of Mathematics, Code MA/Nd

Monterey, CA 93943 USA

Tel: +1 408 656 2235 FAX: +1 408 656 2355

E-mail: bneta@moon.math.nps.navy.mil (Internet)

25 November 2011 Version 1.09

Abstract

This bibliography records publications of Beny Neta.

Title word cross-reference

y′′ = f(t, y) [CRN86].

Adaptive [Net84a, NN87b]. Advection [NW86a, NW86b]. Algorithms [CKMN95, DNE94, NM92]. Analysis [Net81b, Net81c, NV82a, NV82b, NN87a, ND88, NN89, Net89a, NW90, NN90, NN91, Net92a]. Analytic [NDOB94]. Applications [Net92b]. Approximation [NF77, Net77, Net78, NT93a, NT93b]. Approximations [NW89]. Audio [CKMN95, NM92].

Backward [Net86a]. Barotropic [SWN93]. Based [NF84, Net86a]. Benefit [NM93].

Best [Net80b]. Best-Possible [Net80b]. Body [NIL88].

Class [CN84, NL83, Net85]. Cluster [NDOB94, SSN94]. Combat [CKMN95, NM93]. Comparative [WN89]. Computers [NT85]. Computing [ND92]. Confluent [GN89]. Conservative [LN92, NIL88, NL93]. Constants [Net80b]. Convergence [NKN90, NV81b, NV82a, NV82b]. Coordinates [NN90]. Correctors [Net85]. Cowell [Net84b]. Cueing [NM93].

Damped [Net86b, Net89b]. Decomposition [NO90]. Depth [SWN93]. Derivatives [Net80b]. Detection [CKMN95, NM92]. Determination [Net80b, Net92b]. Determining [NV84]. Development [NV81b]. Difference [NV81a, NW89, NV83]. Differences [NI85a, NI85b, NW87, Net92a, WN89]. Differential [CN86, LNK91, LN91, LNG92, Net82a, Net82b, NL83, NF84, Net84b, NI85a, NI85b, Net85, NSvdH85, NO90, Net92b, SvdHN86]. Differentiation [Net86a]. Diffusion [NF77, Net77, NV81b]. Dimensional [ND88, NN90]. Discrete [NV81a, NV82a, NV82b, NV83]. Discrete-Ordinate [NV81a, NV82a, NV82b]. discrete-ordinates [NV83]. Discretization [Net88]. Divergence [NW89, NW90]. Domain [NO90].

Effect [Net88]. Element [NF77, Net77, Net78, NW86a, NW86b, NW89, NW90, Net92a, NT93a, NT93b]. Elements [NI85a, NI85b, NW87, WN89]. Equation [Net82a, Net82b, NW86a, NW86b, NN87b]. Equations [CN86, LN91, Net79, NV81b, Net81d, NV81a, Net81a, NV82a, NV82b, NL83, Net83, Net84a, NF84, Net84b, NV84, NI85a, NI85b, Net85, NSvdH85, Net87, Net88, NN89, Net89a, NW89, NW90, NN90, NO90, NN91, Net92a, NT93a, NT93b, NL93, SvdHN86, LN92, NV83, NN87a, ND88]. Error [Net81b, Net81c]. Evaluation [GN89].

Factorization [NT85]. Families [CN86, NF84, Net86a]. Family [Net79, Net81d, Net83]. Finite [NKN90, NF77, Net77, Net78, NV81a, NI85a, NI85b, NW86a, NW86b, NW87, NW89, NW90, Net92a, NT93a, NT93b, WN89, NV83]. Finite-Difference [NV81a, NV83]. Flow [NW87]. Fluid [NWH86a, NWH86b, NWH86c, NW87]. Form [NW89, NW90]. Formulations [NW86a, NW86b]. Fortran [GN89]. Fourth [CN86, CRN86, NV81a, NV83]. fourth-order [NV83]. Fredholm [Net84a, NN87b]. Function [NN87a, ND88]. Functions [GN89, VN82, VN83].

Group [NW89].

Higher

[Net83, Net84b, NV84, VN82, VN83].

Hybrid [NL83, Net84b, Net85]. Hypercube [LNK91, LNG92, NK89, Net92b, NDOB94]. Hypergeometric [GN89].

II [Net84a]. Implicit [CN84]. Improve [ND92]. Inequalities [Net80b, Net80a]. Influences [SWN93]. Initial [CN84, LNG92, NK89, Net92b, Net94]. Initial-Value [CN84]. Inner [NKN90]. Integral [Net84a, NN87b]. Integrals [Net80b]. Integro [Net82a, Net82b, NI85a, NI85b]. Integro-Differential [Net82a, Net82b, NI85a, NI85b]. INTEL [LNK91, LNG92, Net92b]. Interfaces [Net81b, Net81c]. Involving [Net80b]. Iterations [NKN90]. Iterative [Net81a].

Kelvin [SWN93]. Kernels [Net84a, NN87b]. Kind [Net84a, NN87b].

Lag [CRN86]. Linear [LNK91, NSvdH85, NK89, SvdHN86, SWN93]. Linearly [CN84]. LU [NT85].

MasPar [NT93a, NT93b]. Mathematical [DNE94]. May [ND92]. Method [Net81b, Net81c, NV81a, Net81a, Net84a, NV84, NN87b, VN82, VN83, NV83]. Methods [CN84, CN86, CRN86, Net79, NV81b, Net81d, NV82a, NV82b, NL83, Net83, NF84, Net84b, NSvdH85, Net86a, Net86b, Net89b, NW90, NO90, SvdHN86]. Model [NWH86a, NWH86b, NWH86c, PND93a, PND93b]. Motion [PND93a, PND93b]. Multiple [VN82, VN83]. Multipoint [Net81d]. Multistep [NSvdH85, SvdHN86].

Naval [PND93a, PND93b]. Nonisolated [NV84]. Nonlinear [NF77, Net77, Net78, Net79, Net81d, Net81a, Net82a, Net82b, NV84, NI85a, NI85b, VN82, VN83]. Note [Net81b, Net81c]. Numerical [Net82a, Net82b, Net84a, NN87b].

Orbit [Net92b, NDOB94, SSN94]. Order [CN84, CN86, CRN86, Net79, NV81a, NV82a, NV82b, NL83, Net83, Net84b, NV84, Net85, NSvdH85, SvdHN86, VN82, VN83, NV83]. Ordinary [LNK91, LN91, LNG92, NF84, Net84b, Net92b]. Ordinate [NV81a, NV82a, NV82b]. ordinates [NV83]. Oscillatory [Net86b, Net89b]. Outer [NKN90].

P [CN84, CN86, CRN86]. P-Stable [CN84, CN86, CRN86]. Parabolic [Net78]. Parallel [LN91, NT85, ND92, NL93, Net94, SSN94, LN92]. Parallelization [PND93a, PND93b]. Part [Net84a, NN87b]. Partial [NO90]. Performance [NDOB94]. Periodic [NSvdH85, SvdHN86]. Phase [CRN86, NW86a, NW86b, NW89]. Phase-Lag [CRN86]. Poincaré [SWN93]. Polynomials [NF84, Net86a]. Possible [Net80b]. Prediction [SSN94, WN89]. Predictors [Net85]. Problem [NF77, Net77, Net78, NIL88]. Problems [CN84, LNG92, Net86b, NW87, NK89, Net89b, Net92b, Net94]. Propagator [NDOB94].

Regular [NN87b]. Review [Net92a]. Rossby [NW89].

Satellite [DNE94, PND93a, PND93b, SSN94]. Scheme [LN92, NN87a, NIL88, NN89, Net89a, NN90, NN91, NL93]. Schemes [Net87, ND88]. Second [CN84, CN86, NL83, Net84a, Net85, NSvdH85, NN87b, SvdHN86]. Second-Order [NL83]. Semianalytic [DNE94]. Several [Net87]. Shallow [LN92, NWH86a, NWH86b, NWH86c, NN87a, Net88, ND88, NN89, Net89a, NW89, NW90, NN90, NN91, Net92a, NT93a, NT93b, NL93]. Shallow-Water [NN87a, ND88, NN89]. Simulation [NM93]. Simulations [CKMN95]. Single [CN84]. Single-Step [CN84]. Singular [Net84a]. Six [CRN86]. Sixth [Net79, NV82a, NV82b]. Sixth-Order [Net79, NV82a, NV82b]. Slab [NV81a, NV82a, NV82b, NV83]. Software [LN91, LN92]. Solution [LNK91, LN91, LNG92, Net81a, Net82a, Net82b, Net84a, Net86b, NN87b, Net88, NK89, Net89b, NW90, Net92b, Net94]. Solutions [NV84, NSvdH85, SvdHN86]. Solving [NV81b, NV81a, NV82a, NV82b, Net83, NI85a, NI85b, Net85, Net87, NIL88, NO90, NL93, NV83]. Sound [NM93]. Source [NKN90]. Space [ND92, PND93a, PND93b]. Spatial [Net88]. Special [NL83, Net85, Net86b, Net89b]. Spectral [Net81b, Net81c]. Speed [NW86a, NW86b, NW89]. Sphere [NN91]. Spherical [NN90]. SST [DNE94]. Stability [NW86a, NW86b]. Stable [CN84, CN86, CRN86]. State [Net88]. Steady [Net88]. Steady-State [Net88]. Step [CN84, CN86, CRN86]. Störmer [Net84b]. Studies [NWH86a, NWH86b, NWH86c]. Study [WN89]. Subroutines [GN89]. Surveillance [ND92, PND93a, PND93b]. Symmetric [NSvdH85, SvdHN86]. Synthesis [NV81b, Net81b, Net81c]. System [NV84]. Systems [LNK91, LN91, Net81a].

Termination [NKN90]. Theory [DNE94, NV81b]. Topography [NWH86a, NWH86b, NWH86c]. Transfer [NN87a, ND88]. Transport [NV81a, NV82a, NV82b, NV83]. Trigonometric [NF84, Net86a]. Turkel [NN90, NN91, NN87a, NN89, Net89a].

Turkel-Zwas [NN87a, NN89, Net89a]. Two [CN86, CRN86, NIL88, ND88, NN90]. Two-Step [CN86, CRN86].

Using [SSN94].

Value [CN84, LNG92, NK89, Net92b, Net94]. Variations [SWN93]. Various [NW86a, NW86b, ND88]. Velocities [NW89]. versus [NI85a, NI85b, NW87]. Vorticity [NW89, NW90]. Vorticity-Divergence [NW89, NW90].

Water [LN92, NWH86a, NWH86b, NWH86c, NN87a, Net88, ND88, NN89, Net89a, NW89, NW90, NN90, NN91, Net92a, NT93a, NT93b, NL93]. Wave [NW89]. Waves [SWN93]. Weather [WN89]. whose [Net86b, Net89b]. Workstation [NDOB94, SSN94].

Zeros [VN82, VN83]. Zwas [NN90, NN91, NN87a, NN89, Net89a].

References

Cersovsky:1995:ADA

[CKMN95] D. Cersovsky, E. Kleinschmidt, B. Mansager, and B. Neta. Audio detection algorithms in combat simulations. International J. Mathematical and Computer Modelling, 1995. accepted for publication.

Chawla:1984:NCL

[CN84] M. M. Chawla and B. Neta. A new class of linearly implicit singlestep P-stable methods for second order initial-value problems. Technical Report INTT-28, Institute for Numerical Transport Theory, Texas Tech University, Lubbock, TX, 1984.

Chawla:1986:FTS

[CN86] M. M. Chawla and B. Neta. Families of two-step fourth order Pstable methods for second order differential equations. Journal of Comp. and Appl. Math., 15:213– 223, 1986.

Chawla:1986:TSF

[CRN86] M. M. Chawla, P. S. Rao, and B. Neta. Two-step fourth order P-stable methods with phase-lag of order six for y′′ = f(t, y). Journal of Comp. and Appl. Math., 16: 233–236, 1986.

Danielson:1994:SST

[DNE94] D. A. Danielson, B. Neta, and L. W. Early. Semianalytic satellite theory (SST): Mathematical algorithms. Technical Report NPS-MA-94-001, Department of Mathematics, Naval Postgraduate School, Monterey CA 93943, USA, 1994.

Gragg:1989:FSE

[GN89] W. Gragg and B. Neta. Fortran subroutines for the evaluation of the confluent hypergeometric functions. Technical Report NPS-MA-89-014, Department of Mathematics, Naval Postgraduate School, Monterey CA 93943, USA, 1989. Software available URL http://math.nps.navy.mil.

Lustman:1991:SPS

[LN91] L. Lustman and B. Neta. Software for the parallel solution of systems of ordinary differential equations. Technical Report NPS-MA-91-009, Department of Mathematics, Naval Postgraduate School, Monterey CA 93943, USA, 1991. Software available URL http://math.nps.navy.mil.

Lustman:1992:SPC

[LN92] L. Lustman and B. Neta. Software for the parallel conservative scheme for the shallow water equations. Technical Report NPS-MA-92-004, Department of Mathematics, Naval Postgraduate School, Monterey CA 93943, USA, 1992. Software available URL http://math.nps.navy.mil.

Lustman:1992:SOD

[LNG92] L. Lustman, B. Neta, and W. Gragg. Solution of ordinary differential initial value problems on an INTEL hypercube. Computers and Mathematics with Applications, 23:65–72, 1992.

Lustman:1991:SLO

[LNK91] L. Lustman, B. Neta, and C. P.

Katti. Solution of linear ordinary differential systems on an INTEL hypercube. SIAM J. Scient. Stat. Comput., 12:1480–1485, 1991.

Neta:1988:TFA

[ND88] B. Neta and C. L. DeVito. The transfer function analysis of various schemes for the two dimensional shallow-water equations. Computers and Mathematics with Applications, 16:111–137, 1988.

Neta:1992:PCM

[ND92] B. Neta and D. A. Danielson. Parallel computing may improve space surveillance. ONR Naval Research Review, 44:54, 1992.

Neta:1994:PAO

[NDOB94] B. Neta, D. A. Danielson, S. Ostrom, and S. K. Brewer. Performance of analytic orbit propagator on a hypercube and a workstation cluster. In AIAA-AAS Astrodynamics Program, pages 10–34. American Institute of Aeronautics and Astronautics, Aerospace Center, 370 L’Enfant Promenade, SW, Washington, DC 20024-2518, August 1994. Paper number 94-3706.

Neta:1977:FEAb

[Net77] B. Neta. Finite Element Approximation of a Nonlinear Diffusion Problem. PhD thesis, Carnegie-Mellon University, Schenely Park, Pittsburgh, PA 15213, July 1977.

Neta:1978:FEA

[Net78] B. Neta. Finite element approximation of a nonlinear parabolic problem. Computers and Mathematics with Applications, 4:247– 255, 1978.

Neta:1979:SOF

[Net79] B. Neta. A sixth-order family of methods for nonlinear equations. International J. Computer Mathematics, 7:157–161, 1979.

Neta:1980:I

[Net80a] B. Neta. On 3 inequalities. Computers and Mathematics with Applications, 6:301–304, 1980.

Neta:1980:DBP

[Net80b] B. Neta. On determination of best-possible constants in integrals inequalities involving derivatives. Math. Comp., 35:1191– 1194, 1980.

Neta:1981:NIM

[Net81a] B. Neta. A new iterative method for the solution of systems of nonlinear equations. In Z. Ziegler, editor, Approximation Theory and Applications, pages 249–263. Academic Press, New York, NY, USA, 1981. ISBN 0-12-780650-4.

Neta:1981:NEAa

[Net81b] B. Neta. Note on error analysis for the spectral synthesis method with interfaces. Zeitschrift fur Angewandte Mathematik und Physik, 32:603–608, 1981.

Neta:1981:NEAb

[Net81c] B. Neta. Note on error analysis for the spectral synthesis method with interfaces. Technical Report INTT-12, Institute for Numerical Transport Theory, Texas Tech University, Lubbock, TX, 1981.

Neta:1981:FMM

[Net81d] B. Neta. On a family of multipoint methods for nonlinear equations. International J. Computer Mathematics, 9:353–361, 1981.

Neta:1982:NSNa

[Net82a] B. Neta. Numerical solution of a nonlinear integro-differential equation. J. Math. Anal. Appl., 89:598–611, 1982.

Neta:1982:NSNb

[Net82b] B. Neta. Numerical solution of a nonlinear integro-differential equation. Technical Report INTT-21, Institute for Numerical Transport Theory, Texas Tech University, Lubbock, TX, 1982.

Neta:1983:NFH

[Net83] B. Neta. A new family of higher order methods for solving equations. International J. Computer Mathematics, 14:191–195, 1983.

Neta:1984:AMN

[Net84a] B. Neta. Adaptive method for the numerical solution of Fredholm integral equations of the second kind. part II: Singular kernels. In A. Gerasoulis and R. Vichnevetsky, editors, Numerical Solution of Singular Integral Equations, pages 249–263. IMACS, Department of Computer Science, Rutgers University, New Brunswick, NJ, 1984.

Neta:1984:HOH

[Net84b] B. Neta. Higher order hybrid Störmer–Cowell methods for ordinary differential equations. Congressus Numerantium, 42:251– 264, 1984.

Neta:1985:HPC

[Net85] B. Neta. Hybrid predictors and correctors for solving a special class of second order differential equations. Congressus Numerantium, 46:241–248, 1985.

Neta:1986:FBD

[Net86a] B. Neta. Families of backward differentiation methods based on trigonometric polynomials. International J. Computer Mathematics, 20:67–75, 1986.

Neta:1986:SMP

[Net86b] B. Neta. Special methods for problems whose oscillatory solution is damped. Technical Report NPS-53-86-012, Department of Mathematics, Naval Postgraduate School, Monterey CA 93943, USA, 1986.

Neta:1987:SNS

[Net87] B. Neta. Several new schemes for solving equations. International J. Computer Mathematics, 23:265– 282, 1987.

Neta:1988:ESD

[Net88] B. Neta. The effect of spatial discretization on the steadystate solution of the shallow water equations. In H. Niki and M. Kawahara, editors, Computational Methods in Flow Analysis, volume 2, pages 1041– 1048. Okayama, Japan, September 1988.

Neta:1989:ATZb

[Net89a] B. Neta. Analysis of the Turkel-Zwas scheme for the shallow water equations. In W. F. Ames, editor, Numerical and Applied Mathematics, volume 1(3), pages 257– 264. IMACS, Department of Computer Science, Rutgers University, New Brunswick, NJ, 1989.

Neta:1989:SMP

[Net89b] B. Neta. Special methods for problems whose oscillatory solution is damped. Applied Mathematics and Computation, 31:161– 169, 1989.

Neta:1992:AFE

[Net92a] B. Neta. Analysis of finite element and finite differences for shallow water equations: a review. Mathematics and Computers in Simulation, 34:141–161, 1992.

Neta:1992:SOD

[Net92b] B. Neta. Solution of ordinary differential initial value problems on an INTEL hypercube with applications to orbit determination. In A. J. Coster and K. P. Schwan, editors, 1992 Space Surveillance Workshop, volume 1, pages 205– 208. M.I.T. Lincoln Laboratory, Lexington, MA, April 1992.

Neta:1994:PSI

[Net94] B. Neta. Parallel solution of initial value problems. In D. Bainov, editor, Fourth International Colloquium on Differential Equations. Plovdiv, Bulgaria, August 1994.

Neta:1977:FEAa

[NF77] B. Neta and G. J. Fix. Finite element approximation of a nonlinear diffusion problem. Computers and Mathematics with Applications, 3:287–298, 1977.

Neta:1984:FMO

[NF84] B. Neta and C. H. Ford. Families of methods for ordinary differential equations based on trigonometric polynomials. Journal of Comp. and Appl. Math., 10:33–38, 1984.

Neta:1985:FDVa

[NI85a] B. Neta and J. O. Igwe. Finite differences versus finite elements for solving nonlinear integrodifferential equations. J. Math. Anal. Appl., 112:607–618, 1985.

Neta:1985:FDVb

[NI85b] B. Neta and J. O. Igwe. Finite differences versus finite elements for solving nonlinear integrodifferential equations. J. Math. Anal. Appl., 112:607–618, 1985.

Neta:1988:NCS

[NIL88] B. Neta and Y. Ilan-Lipowski. A new conservative scheme for solving the two body problem. Technical Report NPS-MA-88-009, Department of Mathematics, Naval Postgraduate School, Monterey CA 93943, USA, 1988.

Neta:1989:SLI

[NK89] B. Neta and C. P. Katti. Solution of linear initial value problems on a hypercube. Technical Report NPS-MA-89-001, Department of Mathematics, Naval Postgraduate School, Monterey CA 93943, USA, 1989.

Nelson:1990:CIO

[NKN90] P. Nelson, C. P. Katti, and B. Neta. Convergence of inner– outer source iterations with finite termination of the inner iterations. J. of Integral Equations and Applications, 2:147–174, 1990.

Neta:1983:HMS

[NL83] B. Neta and S. C. Lee. Hybrid methods for a special class of second-order differential equations. Congressus Numerantium, 38:203–225, 1983.

Neta:1993:PCS

[NL93] B. Neta and L. Lustman. Parallel conservative scheme for solving the shallow water equations. Monthly Weather Review, 121: 305–309, 1993.

Neta:1992:ADA

[NM92] B. Neta and B. Mansager. Audio detection algorithms. Technical Report NPS-MA-92-010, Department of Mathematics, Naval Postgraduate School, Monterey CA 93943, USA, 1992.

Neta:1993:BSC

[NM93] B. Neta and B. Mansager. Benefit of sound cueing in combat simulation. Technical Report NPS-MA-94-003, Department of Mathematics, Naval Postgraduate School, Monterey CA 93943, USA, 1993.

Neta:1987:TFA

[NN87a] B. Neta and I. M. Navon. The transfer function analysis of the Turkel-Zwas scheme for the shallow-water equations. Technical Report FSU-SCRI-87-30, SCRI, Florida State University, Tallahassee, FL 32306, 1987.

Neta:1987:AMN

[NN87b] B. Neta and P. Nelson. An adaptive method for the numerical solution of a Fredholm integral equation of the second kind. part I: Regular kernels. Applied Mathematics and Computation, 21:171– 184, 1987.

Neta:1989:ATZa

[NN89] B. Neta and I. M. Navon. Analysis of the Turkel-Zwas scheme for the shallow-water equations. J. Computational Physics, 81:277– 299, 1989.

Neta:1990:ATZ

[NN90] B. Neta and J. Yu I. M. Navon.

Analysis of the turkel zwas scheme for the two dimensional shallow water equations in spherical coordinates. Technical Report FSU-SCRI-90-91, SCRI, Florida State University, Tallahassee, FL 32306, 1990.

Neta:1991:ATZ

[NN91] B. Neta and I. M. Navon. Analysis of the turkel zwas scheme for the shallow water equations on the sphere. In R. Vichnevetsky and J. J. H. Miller, editors, 13th IMACS World Congress on Scientific Computation, volume 1, pages 305–306. Dublin, Ireland, July 1991.

Neta:1990:DDM

[NO90] B. Neta and N. Okamoto. On domain decomposition methods for solving partial differential equations. Technical Report NPS-MA-90-004, Department of Mathematics, Naval Postgraduate School, Monterey CA 93943, USA, 1990.

Neta:1985:SLM

[NSvdH85] B. Neta, B. P. Sommeijer, and P. J. van der Houwen. Symmetric linear multistep methods for second order differential equations with periodic solutions. Technical Report NM-R8501, Center for Mathematics and Computer Science, Amsterdam, The Netherlands, 1985.

Neta:1985:LFP

[NT85] B. Neta and H. M. Tai. LU factorization on parallel computers. Computers and Mathematics with Applications, 11:573–579, 1985.

Neta:1993:FEAa

[NT93a] B. Neta and R. Thanakij. Finite element approximation of the shallow water equations on the maspar. In M. N. Dhaubhadel, M. S. Engelman, and W. G. Habashi, editors, Advances in Finite Element Analysis in Fluid Dynamics, volume 171 of Fluid Engineering Division (FED), pages 47–52. American Society Mech. Engineers, United Engineering Center, 345 E. 47th St., New York, NY 10017, 1993. ISBN 0-7918-1015-1.

Neta:1993:FEAb

[NT93b] B. Neta and R. Thanakij. Finite element approximation of the shallow water equations on the maspar. Technical Report NPS-MA-93-014, Department of Mathematics, Naval Postgraduate School, Monterey CA 93943, USA, 1993.

Neta:1981:NFO

[NV81a] B. Neta and H. D. Victory. A new fourth order finite-difference method for solving discreteordinate slab transport equations. Technical Report INTT-17, Institute for Numerical Transport Theory, Texas Tech University, Lubbock, TX, 1981.

Neta:1981:DCT

[NV81b] B. Neta and H. D. Victory. On the development of a convergence theory of synthesis methods for solving diffusion equations. Progress in Nuclear Energy, 8:283–293, 1981.

Neta:1982:CASa

[NV82a] B. Neta and H. D. Victory.

The convergence analysis for sixth-order methods for solving discrete-ordinate slab transport equations. Numerical Functional Analysis and Optimization, 5:85– 126, 1982.

Neta:1982:CASb

[NV82b] B. Neta and H. D. Victory.

The convergence analysis for sixth-order methods for solving discrete-ordinate slab transport equations. Technical Report INTT-20, Institute for Numerical Transport Theory, Texas Tech University, Lubbock, TX, 1982.

Neta:1983:NFO

[NV83] Beny Neta and H. D. Victory, Jr. A new fourth-order finitedifference method for solving discrete-ordinates slab transport equations. SIAM Journal on Numerical Analysis, 20(1):94–105, February 1983. CODEN SJ-NAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Neta:1984:HOM

[NV84] B. Neta and H. D. Victory. A higher order method for determining nonisolated solutions of a system of nonlinear equations. Computing, 32:163–166, 1984. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Neta:1986:SPSa

[NW86a] B. Neta and R. T. Williams. Stability and phase speed for various finite element formulations of the advection equation. Computers and Fluids, 14:393–410, 1986.

Neta:1986:SPSb

[NW86b] B. Neta and R. T. Williams. Stability and phase speed for various finite element formulations of the advection equation. Computers and Fluids, 14:393–410, 1986.

Neta:1987:FEV

[NW87] B. Neta and R. T. Williams. Finite elements versus finite differences for fluid flow problems. In G. N. Pande and J. Middleton, editors, Transient Dynamic Anal. and Constitutive Laws for Engnrg Materials, pages T11, 1–8. Martinus Nijhoff Pub., Dordrecht, The Netherlands, 1987.

Neta:1989:RWP

[NW89] B. Neta and R. T. Williams.

Rossby wave phase speed and group velocities for finite element and finite difference approximations to the vorticity-divergence form of the shallow water equations. Monthly Weather Review, 117:1439–1457, 1989.

Neta:1990:AFE

[NW90] B. Neta and R. T. Williams. Analysis of finite element methods for the solution of the vorticitydivergence form of the shallow water equations. In M. Yasuhara, H. Daiguji, and K. Oshima, editors, Numerical Methods in Fluid Dynamics, pages 402– 407. Nagoya, Japan, August 28-31, 1989 1990.

Neta:1986:SSWa

[NWH86a] B. Neta, R. T. Williams, and D. E.

Hinsman. Studies in a shallow water fluid model with topography. In R. Vichnevetsky and J. Vignes, editors, Numerical Mathematics and Applications, pages 347–354, 1986.

Neta:1986:SSWb

[NWH86b] B. Neta, R. T. Williams, and D. E.

Hinsman. Studies in a shallow water fluid model with topography. In R. Vichnevetsky and J. Vignes, editors, Numerical Mathematics and Applications, pages 347–354, 1986. ISBN 0-12-780650-

Neta:1986:SSWc

[NWH86c] B. Neta, R. T. Williams, and D. E. Hinsman. Studies in a shallow water fluid model with topography. In R. Vichnevetsky and J. Vignes, editors, Numerical Mathematics and Applications, pages 347–354, 1986. ISBN 0-12-780650-4.

Phipps:1993:PNSa

[PND93a] W. E. Phipps, B. Neta, and D. A. Danielson. Parallelization of the naval space surveillance satellite motion model. J. Astronautical Sciences, 41:207–216, 1993.

Phipps:1993:PNSb

[PND93b] W. E. Phipps, B. Neta, and D. A. Danielson. Parallelization of the naval space surveillance satellite motion model. In R. W. Miller and R. Sridharan, editors, 1993 Space Surveillance Workshop, volume 2, pages 71–79. M.I.T. Lincoln Laboratory, Lexington, MA, April 1993.

Stone:1994:PSO

[SSN94] L. C. Stone, S. B. Shukla, and B. Neta. Parallel satellite orbit prediction using a workstation cluster. International J. Computer and Mathematics with Applications, 28(8):1–8, 1994.

Sommeijer:1986:SLM

[SvdHN86] B. P. Sommeijer, P. J. van der Houwen, and B. Neta. Symmetric linear multistep methods for second order differential equations with periodic solutions. Appl. Num. Math., 2:69–77, 1986.

Staniforth:1993:ILD

[SWN93] A. Staniforth, R. T. Williams, and B. Neta. Influences of linear depth variations on barotropic Kelvin and Poincaré waves. J. of the Atmospheric Sciences, 50:929– 940, 1993.

Victory:1982:HOM

[VN82] H. D. Victory and B. Neta. A higher order method for multiple zeros of nonlinear functions. Technical Report INTT-19, Institute for Numerical Transport Theory, Texas Tech University, Lubbock, TX, 1982.

Victory:1983:HOM

[VN83] H. D. Victory and B. Neta. A higher order method for multiple zeros of nonlinear functions. International J. Computer Mathematics, 12:329–335, 1983.

Williams:1989:CSF

[WN89] R. T. Williams and B. Neta. A comparative study of finite elements and finite differences for weather prediction. In R. Gruber, J. Periaux, and R. P. Shaw, editors, Fifth International Symposium on Numerical Methods in Engineering, volume 1, pages 659–669. Lausanne, Switzerland, September 1989.